What are the quantum numbers for the electrons in the lithium atom (Z =

3) when the atom is in its ground state?

and do the Same question as 5 for Sodium (Z = 11).


1
Expert's answer
2021-12-20T10:25:56-0500

The electron configuration of lithium is 1s22s1

The quantum number sets are:

For two 1s-electrons: (1, 0, 0, "\frac{1}{2}"), (1, 0, 0, -"\frac{1}{2}")

For 2s-electron: (2, 0, 0, "\frac{1}{2}") or (2, 0, 0, -"\frac{1}{2}")


The electron configuration of sodium is 1s22s22p63s1

The quantum number sets are:

For two 1s-electrons: (1, 0, 0, "\frac{1}{2}"), (1, 0, 0, -"\frac{1}{2}")

For two 2s-electrons: (2, 0, 0, "\frac{1}{2}"), (2, 0, 0, -"\frac{1}{2}")

For six 2p-electrons: (3, 1, -1, "\frac{1}{2}"), (3, 1, -1, -"\frac{1}{2}"), (3, 1, 0, "\frac{1}{2}"), (3, 1, 0, -"\frac{1}{2}"), (3, 1, 1, "\frac{1}{2}"), (3, 1, 1, -"\frac{1}{2}")

For 3s-electron: (3, 0, 0, "\frac{1}{2}") or (3, 0, 0, -"\frac{1}{2}")


All quantum number sets are written in the form (n, l, ml, ms).